OpenStudy (anonymous):

Write the equation of the horizontal or vertical line: (-4,5) and m = 0.

OpenStudy (anonymous):

y-y1=m(x-x1) here x1= -4, y1=5 therefore y-5=0(x+4) i.e. y-5 =0 i.e. y=5 Hence the required eq of line is y=5